A geyser shoots water directly upward to a height of 25.4 metres above the ground.

What is the initial velocity of the water, in m/s? (Take the acceleration due to gravity to be g = 9.81 m/s2 .)
Answer #1

height, h= 25.4 m

final veocity, v=0 m/s

Initial Velocity, u= ?

acceleration due to gravity, g= - 9.81 m/s2 (moving against the gravity)

v2=u2+2*g*h

0=u2+2*(-9.81)*25.4

u2 = 2*9.81*25.4

u=22.32m/s